Two consecutive even numbers have a sum of 460. What is the odd number that lies between them ? Need to show equation and all work
Answer by mananth(16949)

let the numbers be x & x+2
x+x+2 = 460
2x+2 = 460
2x= 460-2
2x=458
/2
x=229
x+1 is the even number betwen them
229+1 = 230 is the even number